# Fast offline decoding with local message-passing automata

### v1

@@ Abstract
We present a local offline decoder for topological codes that operates according to a parallelized message-passing framework. The decoder works by passing messages between anyons, with the contents of received messages used to move nearby anyons towards one another. We prove the existence of a threshold, and show that in a system of linear size $L$, decoding terminates with an $O((\log L)^{\eta})$ average-case runtime, where $\eta$ is a small constant. For i.i.d Pauli noise, our decoder has $\eta=1$ and a threshold at a noise strength of $p_{c}\approx 7.3\%$.
@@ 1 Introduction
This work develops local CA decoders for topological codes defined by a simple message-passing framework. Our decoders operate by finding a matching between anyons (violated stabilizers) in a parallel, decentralized fashion: under the decoding dynamics, anyons send out signals broadcasting their locations, and use local feedback based on incoming signals to move nearby anyons towards one another. This scheme is naturally translation invariant, can be implemented in continuous time with a strictly local Lindbladian (with each term having support on $O(1)$ sites), and on a system of linear size $L$, requires access to only $O(\log\log L)$ classical control bits per site.
@@
The decoders we develop in this work are designed for offline decoding: given a logical state ${\rho_{\sf log}}$ and a noise channel $\mathcal{E}$, they recover ${\rho_{\sf log}}$ when input $\mathcal{E}({\rho_{\sf log}})$, under the assumptions that stabilizer measurements can be made perfectly, and that the noise is not too strong (sometimes known as the “code capacity scenario”).111“Offline decoding” could also be used in a situation in which a decoder is run on an accumulated spacetime trajectory of (unreliable) syndrome measurements. Our decoders can also be used in this context. This is enough for establishing mixed-state phase equivalence in the sense discussed above, and our work thus provides an explicit construction of a strictly local Lindbladian that quickly connects noisy topologically ordered states to their clean counterparts (our decoders can be regarded as an alternative to the explicit RG-based decoders developed for this purpose in Ref. [20], which lack translation invariance and require nonlocal operations over $\Omega({\rm polylog}(L))$ length scales).
@@
Offline decoders do not by themselves stabilize fault-tolerant memories. This must be done with an online (or “real-time”) decoder, which operates in a setting where noise occurs continuously and measurement outcomes are unreliable. With the notable exception of codes allowing for single-shot error correction [24, 25, 26]—where the assumption of perfect stabilizer measurements is not needed—the speed of an offline decoder is therefore generally not of direct practical significance. More broadly though, understanding how to quickly perform offline decoding is a necessary stepping stone on the path towards designing efficient fault-tolerant quantum memories, and extensions of the present decoders to this setting will be given in future work [27].
@@
Figure 1: Illustration of the message passing architecture in two dimensions. a) A schematic of how messages are produced when the decoder operates synchronously. A site hosting an anyon (green circle) sends out messages which propagate outward from the anyon’s location at a constant speed (with respect to the $\infty$-norm), and increase their value by 1 at each time step. The decoder makes use of this information by moving each anyon in the direction (±$(\pm{\mathbf{\hat{x}}}$ or ±$\pm{\mathbf{\hat{y}}}$) of the received message with the smallest value (to convey directional information, the decoder actually employs four different types of messages; this distinction is not shown in the figure). b) A snapshot of the decoding dynamics in a system of size $L=75$. Darker colors indicate messages with smaller values, and the directions along which each anyon will move at the next time step are indicated by the white arrows. If the message transmission speed was infinite, each anyon would move in the direction of its nearest neighbor. For finite message speeds, it is possible for an anyon to move towards the location where an anyon was annihilated in the recent past (the rightmost anyon in the figure being an example).
@@ 1.1 Summary and overview
This work focuses on the problem of designing parallelized offline CA decoders for topological stabilizer codes with point-like anyons (in any dimension). Our interest is in decoders which perform computations and apply feedback according to translationally invariant spatially local CA update rules.222 Our emphasis on spatial locality perhaps deserves some elaboration. In a strict sense, any offline decoder that operates non-locally on a system-size-dependent scale $R_{\sf nl}(L)$ can be rendered local by adding additional degrees of freedom that communicate instructions to different parts of the lattice before nonlocal operations are applied. Doing so slows down decoding by an amount at least of order $R_{\sf nl}(L)$, and compromises the simplicity of the decoding scheme. Moreover, this construction leads to an extensively long increase in waiting times between the application of nontrivial feedback operations, which causes problems when trying to extend to the setting of real-time decoding. Our interest in this work is therefore in decoders which are “naturally” local, with the ratio between local classical data-processing steps and nontrivial feedback operations being fixed to an $O(L^{0})$ constant. These decoders are built from an array of classical processors, with one processor located at the location of each minimal stabilizer generator. A processor at site $\mathbf{r}$ is restricted to performing the following actions at each time step (temporarily working in a discrete-time setting for simplicity):
@@
    Measuring the eigenvalue of the stabilizer based at site $\mathbf{r}$, obtaining outcome $\sigma_{\mathbf{r}}$.
@@ 1.1.1 Construction
We begin by briefly describing how our decoders are constructed, focusing on the case of the 2D surface code for concreteness. For simplicity, we will first explain their operation in discrete time, where a global synchronizing clock ensures that all sites of the system update simultaneously at each time step.
@@
    Each anyon emits messages in all directions, which propagate outward from its location at a constant speed $v$. Messages come in four different types, with each type propagating along one of the directions ±$\pm{\mathbf{\hat{x}}},\pm{\mathbf{\hat{y}}}$. Message propagation occurs in such a way that by time $r/v$, an anyon broadcasts messages to all sites in an $\infty$-ball of radius $r$ centered on its location (c.f. Fig. 1).
@@
    An operator $\mathcal{F}$ is applied which moves each anyon by one lattice site in the direction (±$\pm{\mathbf{\hat{x}}}$ or ±$\pm{\mathbf{\hat{y}}}$) of the smallest-valued message it receives.
@@
We will see in Sec. 5 that in a system of linear size $L$, we can take messages to disappear after propagating for time $O(\log L)$ without compromising the existence of a threshold (at the expense of getting a slower—albeit still superpolynomial—suppression of the sub-threshold logical error rate). With this rule, the largest value a message can take is accordingly $O(\log L)$, and hence the number of control bits our decoder uses per site scales as $O(\log\log L)$, as mentioned above. The divergence as $L\rightarrow\infty$ is so weak that we will not be concerned with addressing it. [^4] If $L$ is upper bounded by the number of atoms in the observable universe, $\log_{2}\log_{2}L<10$.
In order to fully parallelize error correction, decoding needs to be performed asynchronously, with each site executing operations independently of a global control clock. We will address this issue by extending our decoder to operate in continuous time, where $\mathcal{A},\mathcal{F}$ are performed by an appropriate Lindbladian. The most naive extension—where each site performs the same updates as in the discrete-time setting—is not a priori guaranteed to work, as the behavior of synchronous cellular automata are generically completely altered in this setting. Nevertheless, we develop a desynchronization scheme which provably allows any discrete-time decoder to be extended to a continuous time Lindbladian without changing the value of its threshold or its expected decoding time. This construction is based on the “marching soldiers” rule of Berman and Simon [28] (see also [29]), and operates by using a continuous time system to simulate the computation performed by a discrete time one (assuming the absence of errors in the classical control bits). The details are provided in Sec. 6.
@@ 1.1.2 Decoding performance
The restriction to this class of stochastic noise models is common in fault tolerance proofs (c.f. Refs. [15, 30, 31, 32]). However, this type of noise does not encompass (rather benign) types of coherent errors, e.g. a small unitary rotation performed on all qubits. The techniques used to treat non-Markovian noise in [33] can be used to extend our analysis to general types of coherent noise, but the details are somewhat involved, and will be deferred to future work [27].
Figure 2: Clustering and error correction under the message-passing decoder, illustrated for the case of the 1D repetition code. a) A spacetime schematic illustrating how the decoder corrects errors. The noise in the input state can be recursively organized into clusters of different sizes. The smallest clusters are indicated by the locations where the blue regions intersect the $t=0$ axis. The decoder is guaranteed to eliminate these clusters before they have a chance to merge with one another, and the spacetime support of anyons contained in these small clusters is restricted to the interiors of the blue diamonds. Disregarding these small clusters, the remaining part of the noise is broken up into larger clusters (yellow regions). The decoder similarly eliminates anyons in these regions before they have a chance to merge. This clustering process is continued up to the scale of the largest clusters in the system (here, the red region), with $p_{\sf log}^{(\mathcal{E})},T_{\sf dec}^{(\mathcal{E})}$ determined by how these large clusters are eliminated. b) elimination of clusters in action. A visualization of the decoding dynamics for a system of size $L=512$, initialized on a random bit string. Black and white areas denoting regions where the bits take values of 00 and $1$, respectively.
@@
Consider the message-passing decoder operating on a $D$-dimensional system of linear size $L$, with either periodic or open boundary conditions, and in either discrete or continuous time. If we allow the decoder to store at least $2D\log_{2}(L)$ control bits at each site, then there exist positive $O(1)$ constants $p_{c},\beta,\eta$ such that as long as $p<p_{c}$, then under any $p$-bounded stochastic noise model $\mathcal{E}$ we have
@@
If we require that the decoder store only $O(\log\log L)$ control bits per site, the bound on the logical error rate is loosened to $p_{\sf log}^{(\mathcal{E})}=(p/p_{c})^{\Omega((\log L)^{\gamma})}$ for a positive $O(1)$ constant $\gamma$, while the bound on $T_{\sf dec}^{(\mathcal{E})}$ remains the same.
@@
We do not attempt to give tight bounds on $p_{c}$ or the allowable values of $\beta$ and $\eta$, and content ourselves with the (likely extremely conservative) bounds $p_{c}>1/29^{D},\beta>\log_{13}(2)\approx 0.27,\,\eta<\log_{2}(13)\approx 3.7$. Without additional assumptions on $\mathcal{E}$ it is in fact impossible to improve the bounds on $\beta,\eta$ to the values of $\beta=\eta=1$ suggested by (2), as in Sec. 5.3 we will construct explicit examples of $p$-bounded noise models with $\beta=\log_{6}(5)<1,\,\eta=\log_{5}(6)>1$. These noise models create fractal-like patterns of anyons and involve highly non-local correlations in space, and we conjecture that $\eta=1$ for noise with short-range spatial correlations.
@@
We prove Theorem 1.1 by formalizing the multi-scale approach schematically described above, showing that sufficiently nearby anyons rapidly annihilate, that anyons at larger scales annihilate after the smaller pairs have disappeared, and so on. The proof consists of two parts, the logic of which is illustrated in Fig. 2. The first part proves a “linear erosion” property [34, 19, 35]: all anyons in a sufficiently well-isolated cluster of linear size $\Delta$ annihilate against one another in a time proportional to $\Delta$, even when they (inevitably) receive signals from anyons outside the cluster. The second part uses a variant of a general renormalization-type argument originally due to Gacs [36, 37]—iterations of which have proven useful in many threshold proofs, see e.g. Refs. [38, 12, 39, 16, 40]—which shows that the anyons created by the noise can be grouped into well-isolated clusters on a hierarchy of increasing length scales, with large clusters rapidly becoming rare provided $p$ is sufficiently small. The theorem then follows from a standard percolation-type argument, with the erosion property showing that different clusters never merge with one another under the decoding dynamics (despite surviving for long enough to exchange messages). $p_{\sf log}^{(\mathcal{E})}$ is then fixed by the probability of the noise creating a system-spanning cluster, and $T_{\sf dec}^{(\mathcal{E})}$ by the typical size of the largest cluster.
@@
In Sec. 7 we use Monte Carlo simulations to quantitatively address the performance of our decoder under i.i.d Pauli noise. Our simplest decoding scheme yields a threshold consistent with $p_{c}=1/2$ in 1D (the repetition code) and $p_{c}\approx 7.3\%$ in 2D, and produces scalings of $p_{\sf log}^{(\mathcal{E})}=\exp(\Theta(L))$, $T_{\sf dec}^{(\mathcal{E})}=\Theta(\log L)$ for $p$ not too much smaller than $p_{c}$, consistent with (2) (for very small $p$, the aforementioned fractal-like noise patterns may dominate the scaling of $p_{\sf log}^{(\mathcal{E})}$). We believe that any local decoder must have $T_{\sf dec}^{(\mathcal{E})}=\Omega(\log L)$ (see Sec. 5.3); if this is true, the scaling of $T_{\sf dec}^{(\mathcal{E})}$ is optimal. Beyond increasing the numerical value of $p_{c}$ by a few (no more than 3 [21]) percent, the only improvement would be to reduce the number of control bits per site to $O(1)$ instead of $O(\log\log L)$. While this is an academic point, in App. E we nevertheless give a construction using only $O(1)$ bits per site that we believe (but do not prove) retains a threshold.

### v4

@@ Abstract
We present a local offline decoder for topological codes that operates according to a parallelized message-passing framework. The decoder works by passing messages between anyons, with the contents of received messages used to move nearby anyons towards one another. We prove the existence of a threshold, and show that in a system of linear size $L$, decoding terminates with an $O((\log L)^{\eta})$ average-case runtime, where $\eta$ is a small constant. For the toric code subject to i.i.d Pauli noise, our decoder has $\eta=1$ and a threshold at a noise strength of $p_{c}\approx 7.3\%$.
@@ 1 Introduction
This work develops local CA decoders for topological codes (Abelian or non-Abelian) defined by a simple message-passing framework. Our decoders operate by finding a matching between anyons in a parallel, decentralized fashion: under the decoding dynamics, anyons send out signals broadcasting their locations, and use local feedback based on incoming signals to move nearby anyons towards one another. This scheme is naturally translation invariant, can be implemented in continuous time with a strictly local Lindbladian (with each term having support on $O(1)$ sites), and on a system of linear size $L$, achieving a suppression of the sub-threshold logical error rate which is superpolynomial in $L$ requires access to only $c\log\log L$ classical control bits per site, where $c$ is a small $O(1)$ constant.
@@
The decoders we develop in this work are designed for offline decoding: given a logical state $\rho_{\sf log}$} and a noise channel $\mathcal{E}$, they recover $\rho_{\sf log}$} when input $\mathcal{E}({\rho_{\sf log}})$, under the assumptions that topological charge measurements can be made perfectly, and that the noise is not too strong (sometimes known as the “code capacity scenario”).111“Offline decoding” could also be used in a situation in which a decoder is run on an accumulated spacetime trajectory of (unreliable) syndrome measurements. Our decoders can also be used in this context. This is enough for establishing mixed-state phase equivalence in the sense discussed above, and our work thus provides an explicit construction of a strictly local Lindbladian that quickly connects noisy topologically ordered states to their clean counterparts (our decoders can be regarded as an alternative to the explicit RG-based decoders developed for this purpose in Ref. [20], which lack translation invariance and require nonlocal operations over $\Omega({\rm polylog}(L))$ length scales).
@@
Offline decoders do not by themselves stabilize fault-tolerant memories. This must be done with an online (or “real-time”) decoder, which operates in a setting where noise occurs continuously and measurement outcomes are unreliable. With the notable exception of codes allowing for single-shot error correction [24, 25, 26]—where the assumption of perfect measurements is not needed—the speed of an offline decoder is therefore generally not of direct practical significance. More broadly though, understanding how to quickly perform offline decoding is a necessary stepping stone on the path towards designing efficient fault-tolerant quantum memories, and extensions of the present decoders to this setting are presented in Ref. [27].
@@
Figure 1: Illustration of the message passing architecture in two dimensions. a) A schematic of how messages are produced when the decoder operates synchronously. A site hosting an anyon (green circle) sends out messages which propagate outward from the anyon’s location at a constant speed (with respect to the $\infty$-norm), and increase their value by 1 at each time step. The decoder makes use of this information by moving each anyon in the direction $(\pm{\mathbf{\hat{x}}}$ or $\pm{\mathbf{\hat{y}}}$) of the received message with the smallest value (to convey directional information, the decoder actually employs four different types of messages; this distinction is not shown in the figure). b) A snapshot of the decoding dynamics in a system of size $L=75$. Darker colors indicate messages with smaller values, and the directions along which each anyon will move at the next time step are indicated by the white arrows. If the message transmission speed was infinite, each anyon would move in the direction of its nearest neighbor. For finite message speeds, it is possible for an anyon to move towards the location where an anyon was annihilated in the recent past (the rightmost anyon in the figure being an example).
@@ 1.1 Summary and overview
This work focuses on the problem of designing parallelized offline CA decoders for topological codes with point-like anyons (in any dimension). Our interest is in decoders which perform computations and apply feedback according to translationally invariant spatially local CA update rules. [^2] Our emphasis on spatial locality perhaps deserves some elaboration. In a strict sense, any offline decoder that operates non-locally on a system-size-dependent scale $R_{\sf nl}(L)$ can be rendered local by adding additional degrees of freedom that communicate instructions to different parts of the lattice before nonlocal operations are applied. Doing so slows down decoding by an amount at least of order $R_{\sf nl}(L)$, and compromises the simplicity of the decoding scheme. Moreover, this construction leads to an extensively long increase in waiting times between the application of nontrivial feedback operations, which causes problems when trying to extend to the setting of real-time decoding. Our interest in this work is therefore in decoders which are “naturally” local, with the ratio between local classical data-processing steps and nontrivial feedback operations being fixed to an $O(L^{0})$ constant. These decoders are built from an array of classical processors, with one processor located at the location of each minimal stabilizer generator. A processor at site $\mathbf{r}$ is restricted to performing the following actions at each time step (temporarily working in a discrete-time setting for simplicity):
@@
    Measuring the topological charge at site $\mathbf{r}$, obtaining outcome $\sigma_{\mathbf{r}}$.
@@ 1.1.1 Construction
We begin by briefly describing how our decoders are constructed. For concreteness, we will specialize in most of the paper to the case of the 2D surface code, so that $\sigma_{\mathbf{r}}\in\{\pm 1\}$. The generalization to arbitrary (potentially non-Abelian) topological codes is essentially immediate, and will be explained in sec. 5.4. For simplicity, we will first explain their operation in discrete time, where a global synchronizing clock ensures that all sites of the system update simultaneously at each time step.
@@
    Each anyon emits messages in all directions, which propagate outward from its location at a constant speed $v$. Messages come in four different types, with each type propagating along one of the directions $\pm{\mathbf{\hat{x}}},\pm{\mathbf{\hat{y}}}$. Message propagation occurs in such a way that by time $r/v$, an anyon broadcasts messages to all sites in an $\infty$-ball of radius $r$ centered on its location (c.f. fig. 1).
@@
    An operator $\mathcal{F}$ is applied which moves each anyon by one lattice site in the direction ($\pm{\mathbf{\hat{x}}}$ or $\pm{\mathbf{\hat{y}}}$) of the smallest-valued message it receives.
@@
We will see in sec. 5 that in a system of linear size $L$, we can take messages to disappear after propagating for time $O(\log L)$ without compromising the existence of a threshold (at the expense of getting a slower, albeit still superpolynomia, suppression of the sub-threshold logical error rate). With this rule, the largest value a message can take is accordingly $O(\log L)$, and the number of control bits our decoder uses per site will then be shown to scale as $c\log_{2}\log_{2}L$, where the constant prefactor $c$ will be shown to be no larger than $10$. The divergence as $L\rightarrow\infty$ is so weak that we will not be concerned with addressing it. [^4] If $L$ is upper bounded by the number of atoms in the observable universe, $\log_{2}\log_{2}L<10$.
In order to fully parallelize error correction, decoding needs to be performed asynchronously, with each site executing operations independently of a global control clock. We will address this issue by extending our decoder to operate in continuous time, where $\mathcal{A},\mathcal{F}$ are performed by an appropriate Lindbladian. The most naive extension—where each site performs the same updates as in the discrete-time setting—is not a priori guaranteed to work, as the behavior of synchronous cellular automata are generically completely altered in this setting. Nevertheless, we develop a desynchronization scheme which provably allows any discrete-time decoder to be extended to a continuous time Lindbladian without changing the value of its threshold or its expected decoding time. This construction is based on the “marching soldiers” rule of Berman and Simon [28, 29] (closely related to the restricted solid-on-solid model of surface growth [30]), and operates by using a continuous time system to simulate the computation performed by a discrete time one (assuming the absence of errors in the classical control bits). The details are provided in sec. 6.
@@ 1.1.2 Decoding performance
The restriction to this class of stochastic noise models is common in fault tolerance proofs (c.f. Refs. [15, 31, 32, 33]). However, this type of noise does not encompass (rather benign) types of coherent errors, e.g. a small unitary rotation performed on all qubits. The techniques used to treat non-Markovian noise in [34, 31] can be used to extend our analysis to general types of coherent noise, but the details are somewhat involved and will be omitted.
Figure 2: Clustering and error correction under the message-passing decoder, illustrated for the case of the 1D repetition code. a) A spacetime schematic illustrating how the decoder corrects errors. The noise in the input state can be recursively organized into clusters of different sizes. The smallest clusters are indicated by the locations where the blue regions intersect the $t=0$ axis. The decoder is guaranteed to eliminate these clusters before they have a chance to merge with one another, and the spacetime support of anyons contained in these small clusters is restricted to the interiors of the blue diamonds. Disregarding these small clusters, the remaining part of the noise is broken up into larger clusters (yellow regions). The decoder similarly eliminates anyons in these regions before they have a chance to merge. This clustering process is continued up to the scale of the largest clusters in the system (here, the red region), with $p_{\sf log}^{(\mathcal{E})},T_{\sf dec}^{(\mathcal{E})}$ determined by how these large clusters are eliminated. b) elimination of clusters in action. A visualization of the decoding dynamics for a system of size $L=512$, initialized on a random bit string. Black and white areas denoting regions where the bits take values of 0 and $1$, respectively.
@@
Consider the message-passing decoder operating on any topological code defined on a $D$-dimensional square lattice of linear size $L$, with either periodic or open boundary conditions, and in either discrete or continuous time. If we allow the decoder to store at least $2D\log_{2}(L)$ control bits at each site, then there exist positive $O(1)$ constants $p_{c},\beta,\eta$ such that as long as $p<p_{c}$, then under any $p$-bounded stochastic noise model $\mathcal{E}$ we have
@@
If we require that the decoder store only $O(\log\log L)$ control bits per site, the bound on the logical error rate is loosened to $p_{\sf log}^{(\mathcal{E})}=(p/p_{c})^{\Omega((\log L)^{\gamma})}$ for a positive $O(1)$ constant $\gamma>1$, while the bound on $T_{\sf dec}^{(\mathcal{E})}$ remains the same.
@@
We do not attempt to give tight bounds on $p_{c}$ or the allowable values of $\beta$ and $\eta$, and content ourselves with the (likely extremely conservative) bounds $p_{c}>1/29^{D},\beta>\log_{13}(2)\approx 0.27,\,\eta<\log_{2}(13)\approx 3.7$. Without additional assumptions on $\mathcal{E}$ it is in fact impossible to improve the bounds on $\beta,\eta$ to the values of $\beta=\eta=1$ suggested by (2), as in sec. 5.3 we will construct explicit examples of $p$-bounded noise models with $\beta=\log_{6}(5)<1,\,\eta=\log_{5}(6)>1$. These “Gerrymandered” noise models create fractal-like patterns of anyons and involve highly non-local correlations in space, and we conjecture that $\eta=1$ for noise with short-range spatial correlations.
@@
We prove Theorem 1.1 by formalizing the multi-scale approach schematically described above, showing that sufficiently nearby anyons rapidly annihilate, that anyons at larger scales annihilate after the smaller pairs have disappeared, and so on. The strategy of the proof is standard and consists of two parts, the logic of which are illustrated in fig. 2. The first part proves a “linear erosion” property [35, 19, 36]: all anyons in a sufficiently well-isolated cluster of linear size $\Delta$ annihilate against one another in a time proportional to $\Delta$, even when they (inevitably) receive signals from anyons outside the cluster. The second part uses a variant of a general renormalization-type argument originally due to Gacs [37, 38]—iterations of which have proven useful in many threshold proofs, see e.g. Refs. [39, 12, 40, 16, 41]—which shows that the anyons created by the noise can be grouped into well-isolated clusters on a hierarchy of increasing length scales, with large clusters rapidly becoming rare provided $p$ is sufficiently small. The theorem then follows from a standard percolation-type argument, with the erosion property showing that different clusters never merge with one another under the decoding dynamics, despite surviving for long enough to exchange messages with one another. $p_{\sf log}^{(\mathcal{E})}$ is then fixed by the probability of the noise creating a system-spanning cluster, and $T_{\sf dec}^{(\mathcal{E})}$ by the typical size of the largest cluster.
@@
In sec. 7 we use Monte Carlo simulations to quantitatively address the performance of our decoder for surface codes under i.i.d Pauli noise. Our simplest decoding scheme yields a threshold consistent with $p_{c}=1/2$ in 1D (the repetition code) and $p_{c}\approx 7.3\%$ in 2D, and produces scalings consistent with $p_{\sf log}^{(\mathcal{E})}=\exp(\Theta(L))$, $T_{\sf dec}^{(\mathcal{E})}=\Theta(\log L)$ for $p$ not too much smaller than $p_{c}$, consistent with (2) (for very small $p$, the aforementioned fractal-like noise patterns may dominate the scaling of $p_{\sf log}^{(\mathcal{E})}$). We believe that any local decoder must have $T_{\sf dec}^{(\mathcal{E})}=\Omega(\log L)$ (see sec. 5.3); if this is true, the scaling of $T_{\sf dec}^{(\mathcal{E})}$ is optimal. Beyond increasing the numerical value of $p_{c}$ by a few (no more than 3 [21]) percent, the only improvement would be to reduce the number of control bits per site to $O(1)$ instead of $O(\log\log L)$. While this is an academic point, in app. E we nevertheless give a construction using only $O(1)$ bits per site that we believe (but do not prove) retains a threshold.